OpenStudy (anonymous):

Help Me Please! The mechanical energy of an ideal mass-spring oscillating system starts at 100 joules for a given amplitude. Determine the new mechanical energy of the system. 1) Amplitude is cut in half -> Energy? 2) Amplitude is increased by 30% -> Energy? 3) Amplitude is decreased by 30% -> Energy? I know that energy is proportional to amplitude^2.

rvc (rvc):

Suppose amplitude is doubled then energy will be?

rvc (rvc):

\[E\ \alpha\ a^2\]

rvc (rvc):

now a=2a then E=?

OpenStudy (anonymous):

E would be x4 more

OpenStudy (anonymous):

Which would be 400 Joules, but how do you do the percentages and fractions?

OpenStudy (anonymous):

For example if you do 30% increase in energy I believe I need to take 1.3^2 which is 1.69, so that means it will be an increase in 69% of energy meaning 169 Joules for #2. If I did that correctly.

rvc (rvc):

amplitude is increased by 30% so energy will increase by the square of 30%,am i right?@HelpMe94

OpenStudy (anonymous):

So its just a 9% increase?

rvc (rvc):

hey its the square of 30%

OpenStudy (anonymous):

So if I did #1 that means its (1/2)^2 which is 1/4 so is that just a 25% decrease and #2 with the same 9% decrease.

rvc (rvc):

@amorfide will u please assist us

OpenStudy (amorfide):

me no comprehende physics

OpenStudy (amorfide):

sorry

OpenStudy (anonymous):

I think its more about proportions.

rvc (rvc):

@amorfide i need help with those % thing

OpenStudy (amorfide):

I can not confirm your percentages are correct as I do not know where they came from and such Let me do some research on this

OpenStudy (anonymous):

I am thinking you need to use 1.3^2 not 0.3 because you have increase to 130% of the original value.

OpenStudy (amorfide):

well i cant find anything on it but if part 1 makes senese, then the others should halving it simply decreased it by 50% right?

OpenStudy (amorfide):

so how is increasing by 30% any different, same method so I would assume @rvc is correct

OpenStudy (amorfide):

@rational help us out

rvc (rvc):

if amplitude is increased by 30% then energy is increased by 90%

OpenStudy (anonymous):

i am not sure if i follow...

OpenStudy (anonymous):

E α A^2 Therefore, E=k.A^2 [k is any constant] E=100 J k=100/A^2 when A=A/2 [cut in half] then E=k.A^2  E=100/A^2*(A/2)^2  E=100/4=25 J when A=130/100*A [A increased by 30%] then E=k.A^2  E=100/A^2*(A/2)^2  E=100*(130/100)^2  E=169 J when A=70/100*A [A decreased by 30%] then E=k.A^2  E=100/A^2*(A/2)^2  E=100*(70/100)^2  E=49 J.

OpenStudy (anonymous):

I hope this was what you meant...